OpenStudy (saifoo.khan):

f(x) = -0.6 x^2+1.2 x+1.4 What is the Domain & Range?

OpenStudy (anonymous):

didn't we just do this damned problem?

OpenStudy (saifoo.khan):

And please i want the COMPLETING SQUARE FORM STEPS.

OpenStudy (saifoo.khan):

haha, yes we did.

OpenStudy (anonymous):

so you want all the steps, not the \[-\frac{b}{2a}\] shortcut?

OpenStudy (saifoo.khan):

i have to show it..

OpenStudy (anonymous):

ok fine. here we go. 1) factor out the leading coefficient from the first two terms and get \[f(x)=-.06(x^2-2x)+1.4\]

OpenStudy (anonymous):

2) then take half the middle term, which is -1, square it, get 1, and make sure that you have what you start with because you are not in fact adding 1 but (by the distributive law) subtracting -.6, so you have to add it back \[f(x)=-.6(x^2-2x+1)+1.4+.6\] \[f(x)=-.6(x^2-2x+1)+2\]

OpenStudy (anonymous):

3) write as a perfect square \[f(x)=-.6(x-1)^2+2\]

OpenStudy (saifoo.khan):

THanks! :)

OpenStudy (saifoo.khan):

You are the BOMB.

OpenStudy (anonymous):

4) now the range, max, and vertex are displayed to see vertex \[(1,2)\] max : 2 range \[(-\infty,2)\]

OpenStudy (anonymous):

5) don't ask any more of these damned decimal questions!

OpenStudy (saifoo.khan):

6) Alright sir, that's why i shot that one and made another(easy) one. ;)

OpenStudy (saifoo.khan):

=)

OpenStudy (saifoo.khan):

Where is DOMAIN? :O

OpenStudy (anonymous):

domain of a polynomial unless otherwise specified is all real numbers

OpenStudy (saifoo.khan):

Aint range gonna be: (-infinity, 2]

OpenStudy (anonymous):

what did i write? oh yes, closed, you are right i wrote open. i am wrong and you are right

OpenStudy (saifoo.khan):

lol. thanks again.